OpenStudy (gracygirl):

Simplify (3a^2)^3 (2 and the last 3 are exponents) I just wanted to double check my answer to make sure I'm doing this correct. The answer would be: (6a)^6 Am I correct? Do the parenthesis need to be removed?

OpenStudy (gracygirl):

@sleepyjess @jabez177

OpenStudy (gracygirl):

@pooja195 @mathstudent55

OpenStudy (gracygirl):

@deltaforce209

OpenStudy (deltaforce209):

32a2

OpenStudy (gracygirl):

So...after multiplying the 6 by the exponent 2 it needs to be multiplied by the exponent 6?

OpenStudy (deltaforce209):

thats what i came up with so idk if i am on point correct but im pretty sure it is

563blackghost (563blackghost):

You need to distribute the power of 3 into the equation.... \(\Large{(3^3a^{(2\times 3)})}\)

OpenStudy (deltaforce209):

3a2 • 3 Step 2 : Multiplying exponents : 2.1 31 multiplied by 31 = 3(1 + 1) = 32 Final result : 32a2

OpenStudy (gracygirl):

So...I feel like I'm getting conflicting answers or I just don't see how one plays out right...

OpenStudy (deltaforce209):

which one mine i asked my teacher so it should be right

OpenStudy (gracygirl):

Delta's sounds right...most of the time I just don't take it far enough and it's usually a much bigger number than what I came up with.

OpenStudy (deltaforce209):

thank you

563blackghost (563blackghost):

Its not quite right. You need to distribute the exponent 3 into the equation which having two exponent like in this equation would multiply them....

OpenStudy (gracygirl):

But what would that equal out to?

563blackghost (563blackghost):

Yes it does... \(\LARGE{(3^{3}a^{(2\times 3)})}\) Try simplifying.

OpenStudy (gracygirl):

28a5? Or does the entire thing need to be multiplied by the exponent 5?

OpenStudy (gracygirl):

Wait, 27 not 28.

563blackghost (563blackghost):

Almost there :) 2 by 3 actually equals 6 since we have 3+3 or adding 3 twice...so it simplify to... \(\LARGE{27a^{6}}\)

OpenStudy (gracygirl):

Oh shoot I read the x as a +. So it should be 27a6.

563blackghost (563blackghost):

Yup :)

OpenStudy (gracygirl):

Okay, thank you! Both of you.

563blackghost (563blackghost):

np :)

OpenStudy (deltaforce209):

np